# DomainDemo: a dataset of domain-sharing activities among different demographic groups on Twitter

### v1

@@ Abstract
Social media play a pivotal role in disseminating web content, particularly during elections, yet our understanding of the association between demographic factors and political discourse online remains limited. Here, we introduce a unique dataset, DomainDemo, linking domains shared on Twitter (X) with the demographic characteristics of associated users, including age, gender, race, political affiliation, and geolocation, from 2011 to 2022. This new resource was derived from a panel of over 1.5 million Twitter users matched against their U.S. voter registration records, facilitating a better understanding of a decade of information flows on one of the most prominent social media platforms and trends in political and public discourse among registered U.S. voters from different sociodemographic groups. By aggregating user demographic information onto the domains, we derive five metrics that provide critical insights into over 129,000 websites. In particular, the localness and partisan audience metrics quantify the domains’ geographical reach and ideological orientation, respectively. These metrics show substantial agreement with existing classifications, suggesting the effectiveness and reliability of DomainDemo’s approach.
@@
We release two variants of the domain-sharing statistics dataset: DomainDemo-multivariate and DomainDemo-univariate. The DomainDemo-multivariate variant splits the statistics into buckets defined by domain, age, gender, race, political affiliation, state, and month altogether. In each bucket, we provide the number of shares, the number of unique users sharing the domain, and the Gini index (Gini, 1912) of the sharing count among users. The DomainDemo-univariate variant includes five universes: state, race, gender, age, and political affiliation. Within each universe, we provide the number of shares, the number of unique users sharing the domains, and the Gini index of the sharing count among users for each category (e.g., age groups in the age universe).
@@
The DomainDemo-multivariate variant is similar in format to the Facebook URL dataset shared through Social Science One (King and Persily, 2020; Messing et al., 2020), which includes the number of views, shares, and reactions to various URLs across different demographic groups. However, there are key differences. DomainDemo-multivariate focuses solely on a sample of Twitter users who are registered voters in the U.S., while the Facebook URL dataset includes data from all eligible users on the platform. Additionally, DomainDemo-multivariate only includes the number of shares and unique users at the domain level and does not incorporate any noise.
@@
In addition to the count statistics, we introduce five derived metrics that quantify how the demographics (age, gender, race, political affiliation, and geolocation) of users sharing a particular domain differ from that of the baseline. These metrics have specific interpretations that are useful for various research questions. For instance, the geolocation of the users allows us to measure the domains’ localness, i.e., the extent to which the sharing of a particular domain is geographically localized. Together with the user-sharing behavior data, the localness metric enables researchers to quantify the changing landscape of the local news industry. As a fundamental component of the U.S. democratic process, local news is uniquely positioned to report on local affairs and elections (Martin and McCrain, 2019; Moskowitz, 2021), but faces a declining trend over the years (Hayes and Lawless, 2018; Abernathy, 2018). Similarly, the user party affiliation in DomainDemo allows us to measure the audience partisanship of different domains. This metric can serve as a proxy for the political leaning of the domains, crucial for understanding online political discourse (Bakshy et al., 2015; Robertson et al., 2023).
Our derived metrics demonstrate strong alignment with established measures of localness and political leaning while significantly expanding coverage to over 129,000 domains—over ten times the number of domains in existing datasets. The metrics also uncover subtle variations in sharing patterns that previous binary or one-dimensional categorization schemes could not capture. Furthermore, our dataset enables researchers to analyze how domain-sharing behaviors have evolved across an 11-year period.
@@ Methods
In this section, we describe how we create our dataset.
@@
Our dataset is based on a panel of over 1.5 million Twitter users matched with U.S. voter registration records. We leverage the methodology outlined in previous studies (Grinberg et al., 2019; Shugars et al., 2021; Hughes et al., 2021), utilizing the Twitter Decahose, a 10% random sample of all tweets, to identify 290 million profiles that post content between January 2014 and March 2017. These profiles are then matched with voter records provided by TargetSmart in October 2017, covering all 50 U.S. states and the District of Columbia. A Twitter account is considered a match to a voter record if the full name exactly matches and the individual is the only person with that name in the specified city or state-level geographic area in both datasets. This reliance on full names and disclosed locations helps to eliminate many fake, automated (bot), and organizational accounts.
@@
Matching to voter file records provides access to the geolocation, year of birth, gender, race/ethnicity, and partisanship of the matched Twitter users. For geolocation, we consider only the state of residence as recorded in the voter files. Using the year of birth, we determine the age of users at the time of sharing events and categorize them into the following age groups: “<18,” “18-29,” “30-49,” “50-64,” and “>65.” The category for users younger than 18 years old is included because they might post content before reaching the legal voting age. Gender is a binary measure provided by TargetSmart, which does not capture gender identities beyond the binary framework (Shugars et al., 2024). Race/ethnicity information is inferred by TargetSmart for most states and is categorized as “African-American,” “Asian,” “Hispanic,” “Caucasian,” and “Other.”
@@
TargetSmart provides two measures of partisanship: party registration and inferred partisanship. Party registration information in voter files is self-reported and aligns well with survey self-reporting (Ansolabehere and Hersh, 2017). However, this information is unavailable for 20 states (AL, AR, GA, HI, IL, IN, MI, MN, MO, MS, MT, ND, OH, SC, TN, TX, VA, VT, WA, and WI) in the TargetSmart data, which account for 42.7% of the Twitter users in our panel. When categorizing party registration information, we treat values for users in the 20 aforementioned states as missing. For other states, users registered as “Democrat” and “Republican” are coded accordingly. Due to variations in the classification of independent registered voters by state, we group individuals listed as “Independent,” “No party,” or “Unaffiliated” into a single “Independent” category. Members of minor parties, such as the Green Party and the Libertarian Party, are categorized as “Other.” Based on party registration and other indicators, TargetSmart infers the probability of each person voting Democrat. An evaluation from the Pew Research Center suggests this inferred partisanship is reasonably accurate (Igielnik et al., 2018). We categorize individuals as Republican (0-0.35), Independent (0.35-0.65), and Democrat (0.65-1) using TargetSmart’s recommended thresholds to generate the inferred partisanship. For our data release and analysis, we use inferred partisanship as the primary measure since it covers all users (referred to as “party” hereafter). Additionally, we provide party registration information as a secondary measure (referred to as “party registration” hereafter), as it offers useful insights for certain analyses.
@@
Domain sharing statistics
@@
We collect posts from users in the panel spanning from May 2011 to April 2022. We extract the links shared by these users, expand the shortened links when possible, and identify the corresponding domains (e.g., nytimes.com for The New York Times). This process allows us to determine which user shares which domain and when. Sharing events, as defined in our study, include posting links in original tweets and retweeting or quoting tweets containing links. To reduce noise and the risks of re-identification, we include only domains that are shared by at least 50 unique users throughout the entire period.
@@
We integrate the demographic information of users with their domain-sharing records to construct a comprehensive table. This domain-sharing event table includes the following columns: user_id, domain, age, gender, race, party, party registration, state, and year-month. Each row corresponds to a single sharing event, with users who share the same domains multiple times contributing multiple rows. Due to the presence of user identifiers, we cannot release this detailed table. Instead, we provide aggregate statistics derived from this table, as detailed below.
@@
To produce DomainDemo-multivariate, we group the domain-sharing event table by domain, age, gender, race, party (party registration is excluded here), state, and year-month, and calculate the following statistics: the number of shares, the number of unique users who share the domain, and the Gini index of the sharing count across users. Formally, the Gini index $G$ for a domain is calculated as:
$G$ $=\frac{2}{N^{2}\bar{x}}\sum_{i=1}^{N}\sum_{j=1}^{N}|x_{i}-x_{j}|\text{,}$ (1)
@@
where $N$ is the number of users who share the domain, $x_{i}$ is the number of shares by the $i$th user, and $\bar{x}=\sum_{i=1}^{N}x_{i}/N$ is the average number of shares per user. $G$ ranges from 0 to 1, with 0 indicating equal sharing and values close to 1 indicating that a few users share the domain disproportionately. Note that we set the Gini index to 1 for domains shared by only one user in the bucket. We include the Gini index to help researchers understand the inequality of sharing events across users without releasing detailed information about these users due to privacy concerns.
@@
DomainDemo-univariate is generated by aggregating the sharing events across all demographic dimensions except for the one of interest. For example, the state universe is produced by aggregating the sharing events across all age, gender, race, and party dimensions, resulting in the statistics of sharing events within different states. Similar to DomainDemo-multivariate, in each bucket of DomainDemo-univariate, we provide three statistics: the number of shares, the number of unique users who share the domain, and the Gini index of the sharing count across users.
@@
Demographic dimensions with multiple categories
Consider a demographic dimension with multiple categories indexed by $s$, such as states for geolocation. We first calculate the user distribution across all categories for all domains as the baseline distribution. We then compare the user distribution of individual domains in each category with the baseline distribution to measure the deviation. For geolocation, the deviation can be interpreted as the localness of the domain. This method can be applied to other demographic dimensions with multiple categories, such as age and race.
For a formal definition, let $C_{s}$ represent the number of unique users in category $s$ across all domains, and $F_{s}$ represent the corresponding frequency, where $F_{s}=C_{s}/\sum_{s}C_{s}$. For a domain $\delta$, we calculate the user frequency in category $s$, $F_{\delta,s}=C_{\delta,s}/\sum_{s}C_{\delta,s}$, where $C_{\delta,s}$ represents the number of unique users in category $s$ who share the domain $\delta$. For domains shared by diverse users, the observed distribution $F_{\delta,s}$ should closely align with the baseline distribution $F_{s}$ across different categories. However, deviations from the baseline distribution are expected for domains with a more concentrated audience.
@@
where $F_{\delta,s}\log_{2}(F_{\delta,s}/F_{s})$ measures the discrepancies between the observed sharing patterns and the baseline distribution of domain $\delta$ in category $s$. $\mathcal{L}_{\delta}$ is a non-negative value that is minimized at zero when $F_{\delta,s}$ and $F_{s}$ are identical.
@@
A limitation of $\mathcal{L}_{\delta}$ is that it can only indicate the deviation of a domain’s sharing pattern from the baseline distribution. To reveal which categories are over-represented or under-represented, one needs to further inspect the values of $F_{\delta,s}$ and $F_{s}$.
Demographic dimensions with two categories
Demographic dimensions with two categories are simpler to analyze, allowing us to directly measure the audience leaning of a domain based on the number of users from each category. For instance, in the party dimension, excluding Independent users leaves us with Democrat and Republican users. We can then assess the audience partisanship of a domain by the number of users from each party. This approach is also applicable to the gender dimension.
@@
Formally, the audience leaning score $\mathcal{P}_{\delta}$ of a domain $\delta$ is calculated as follows:
$\mathcal{P}_{\delta}$ $=\frac{\frac{C_{\delta,R}}{C_{R}}-\frac{C_{\delta,L}}{C_{L}}}{\frac{C_{\delta,R}}{C_{R}}+\frac{C_{\delta,L}}{C_{L}}}\text{,}$ (3)
@@
where $C_{\delta,L}$ and $C_{\delta,R}$ represent the number of unique users from the $L$ and $R$ categories who share the domain $\delta$, respectively. $C_{L}$ and $C_{R}$ represent the total number of unique users in the $L$ and $R$ categories who share any domain, respectively. Since a user can share multiple domains, we have $C_{\delta,L}\leq C_{L}\leq\sum_{i\in I}C_{i,L}$ and $C_{\delta,R}\leq C_{R}\leq\sum_{i\in I}C_{i,R}$. $\mathcal{P}_{\delta}$ is a continuous value between $-1$ and $+1$, where $-1$ means the domain $\delta$ is exclusively shared by users in the $L$ category and $+1$ means $\delta$ is exclusively shared by $R$ category users.
@@
Domain sharing statistics
Table 1: Schema of the DomainDemo-multivariate tables.
Column Data type Category Note
domain string key second-level domains of websites, e.g., nytimes.com
state string demographic two-letter aberrations, e.g., MA; including 50 U.S. states and the District of Columbia
@@
time string time year and month in the format “YYYY-MM”
shares integer statistic sharing event count of the domain in the bucket
users integer statistic unique number of users who share the domain in the bucket

### v2

@@ Abstract
Social media play a pivotal role in disseminating web content, particularly during elections, yet our understanding of the association between demographic factors and information sharing online remains limited. Here, we introduce a unique dataset, DomainDemo, linking domains shared on Twitter (X) with the demographic characteristics of associated users, including age, gender, race, political affiliation, and geolocation, from 2011 to 2022. This new resource was derived from a panel of over 1.5 million Twitter users matched against their U.S. voter registration records, facilitating a better understanding of a decade of information flows on one of the most prominent social media platforms and trends in political and public discourse among registered U.S. voters from different sociodemographic groups. By aggregating user demographic information onto the domains, we derive five metrics that provide critical insights into over 129,000 websites. In particular, the localness and partisan audience metrics quantify the domains’ geographical reach and ideological orientation, respectively. These metrics show substantial agreement with existing classifications, suggesting the effectiveness and reliability of DomainDemo’s approach.
@@
We release two versions of the domain-sharing statistics dataset: DomainDemo-multivariate and DomainDemo-univariate. DomainDemo-multivariate splits the statistics into buckets defined by age, gender, race, political affiliation, state, and time altogether. In each bucket, we provide the number of shares, the number of unique users sharing the domain, and the Gini index (Gini, 1936) of the sharing count among users. DomainDemo-univariate includes five universes: state, race, gender, age, and political affiliation. Within each universe, we provide the number of shares, the number of unique users sharing the domains, and the Gini index of the sharing count among users for each category (e.g., age groups in the age universe). In addition to the domain-level statistics, we also provide the distribution of the population in different demographic buckets for both DomainDemo-multivariate and DomainDemo-univariate. These population-level distributions can serve as baselines.
@@
DomainDemo-multivariate is similar in format to the Facebook URL dataset shared through Social Science One (King and Persily, 2020; Messing et al., 2020), which includes the number of views, shares, and reactions to various URLs across different demographic groups. However, there are some key differences. DomainDemo-multivariate focuses solely on a sample of Twitter users who are registered voters in the U.S., while the Facebook URL dataset includes data from all eligible users on the platform. Additionally, DomainDemo-multivariate only includes the number of shares and unique users at the domain level and does not incorporate any noise. However, our dataset provides more detailed demographic information compared to the Facebook URL dataset. Specifically, while the Facebook URL dataset only includes country-level geographic data, our dataset contains state-level geolocation information and includes race information.
@@
In addition to the count statistics, we introduce five derived metrics that quantify how the demographics (age, gender, race, political affiliation, and geolocation) of users sharing a particular domain differ from that of the baseline. These metrics have specific interpretations that are useful for various research questions. For instance, the geolocation of the users allows us to measure the domains’ localness, i.e., the extent to which the sharing of a particular domain is geographically localized. Together with the user-sharing behavior data, the localness metric enables researchers to quantify the changing landscape of the local news industry. As a fundamental component of the U.S. democratic process, local news is uniquely positioned to report on local affairs and elections (Martin and McCrain, 2019; Moskowitz, 2021), but faces a declining trend over the years (Hayes and Lawless, 2018; Abernathy, 2018). Similarly, the user party affiliation in DomainDemo allows us to measure the audience partisanship of different domains. This metric can serve as a proxy for the political leaning of the domains, crucial for understanding online political discourse (Bakshy et al., 2015; Robertson et al., 2023). Our derived metrics demonstrate strong alignment with established measures of localness and political leaning while significantly expanding coverage to over 129,000 domains—over ten times the number of domains in existing datasets. The metrics also uncover subtle variations in sharing patterns that previous binary or one-dimensional categorization schemes could not capture.
Due to the difficulty of obtaining data from social media, especially Twitter, in the post-API era (Murtfeldt et al., 2024), replicating our efforts is challenging. Even if access to Twitter data were to become available in the future, the platform itself has undergone significant changes. These factors make our dataset a unique and valuable contribution to the research community, as it provides a comprehensive view of domain-sharing behaviors across an 11-year period.
@@ Methods
In this section, we describe how we create our dataset and provide case studies to interpret the data.
@@
Our dataset is based on a panel of over 1.5 million registered U.S. voters on Twitter, created by our team in previous work. A pilot version of the panel was first used by Grinberg et al. (2019), then the panel was expanded considerably by  Shugars et al. (2021) and validated by Hughes et al. (2021). To create the panel, we start with the Twitter Decahose, a 10% random sample of all tweets, and identify 290 million accounts that post content between January 2014 and March 2017. We extract the names of the users, either from the Twitter handles or display names, and their location from the account profiles. This information is then matched against voter data provided by TargetSmart in October 2017, covering all 50 U.S. states and the District of Columbia. We compare the full name of each person in the voter file with the names of the Twitter accounts. If the full name has fewer than 10 exact matches, we then examine the location of the Twitter accounts. A Twitter account and voter record pair is accepted only if that is the only person in the specified city or state-level geographic area in both datasets. This reliance on full names and disclosed locations helps to eliminate many fake, automated (bot), and organizational accounts.
The data collection and matching of the Twitter panel were approved by the Northeastern University Institutional Review Board (protocol number: 17-12-13). Following the best practices outlined by Hemphill et al. (2022), we employ data aggregation, anonymization, and access control measures to protect user privacy and minimize the risk of re-identification in our Twitter panel.
@@
Matching to voter file records provides access to the geolocation, year of birth, gender, race/ethnicity, and partisanship of the users in the panel. We use state-level geolocation data from the voter files as our geographic unit of analysis. While we have access to more detailed location information, such as county-level data, releasing this information would risk re-identifying the users due to the low population density of many U.S. counties. State-level granularity offers a good compromise between the usefulness of the data and the privacy of the users. Using the year of birth, we determine the age of users at the time of sharing events and categorize them into the following age groups: “<18,” “18-29,” “30-49,” “50-64,” and “65+.” The category for users younger than 18 years old is included because some states allow 17-year-olds to pre-register to vote, and some users might be younger than 18 at the time of sharing events. Gender is a binary measure provided by TargetSmart, which does not capture gender identities beyond the binary framework (Shugars et al., 2024). Race/ethnicity information is inferred by TargetSmart for most states and is categorized as “African-American,” “Asian,” “Hispanic,” and “Caucasian.” Other race categories with limited representation in the dataset are aggregated into a single “Other” category to minimize re-identification risks.
@@
TargetSmart provides two measures of partisanship: party registration and inferred partisanship. Party registration information in voter files is self-reported and aligns well with survey self-reporting (Ansolabehere and Hersh, 2017). However, this information is unavailable for 20 states (AL, AR, GA, HI, IL, IN, MI, MN, MO, MS, MT, ND, OH, SC, TN, TX, VA, VT, WA, and WI) in the TargetSmart data, which account for 42.7% of the Twitter users in our panel. When categorizing party registration information, we treat values for users in the 20 aforementioned states as missing. For the other 30 states and the District of Columbia, users registered as “Democrat” and “Republican” are coded accordingly. Due to variations in the classification of independent registered voters by state, we group individuals listed as “Independent,” “No party,” or “Unaffiliated” into a single “Independent” category. Members of minor parties, such as the Green Party and the Libertarian Party, are categorized as “Other.”
Based on party registration and other indicators, TargetSmart infers the probability of all individuals in all 50 states and the District of Columbia voting Democrat. We categorize individuals as Republican (0-0.35), Independent (0.35-0.65), and Democrat (0.65-1) using TargetSmart’s recommended thresholds to generate the inferred partisanship. For our data release and analysis, we use inferred partisanship as the primary measure since it covers all users (referred to as “party” hereafter). Additionally, we provide party registration information as a secondary measure (referred to as “party registration” or “partyreg” hereafter), as it conveys a slightly different signal and offers useful insights for certain analyses.
@@
Domain-sharing Statistics
@@
We collect posts from users in the panel spanning from May 2011 to April 2022. We extract the links shared by these users, expand the shortened links when possible, and identify the corresponding domains (e.g., nytimes.com for The New York Times). This process allows us to determine which user shares what domains and when. Sharing events, as defined in our study, include posting links in original tweets and retweeting or quoting tweets containing links. To reduce noise and the risks of re-identification, we include only domains that are shared by at least 50 unique users throughout the entire period.
@@
We integrate the demographic information of users with their domain-sharing records to construct a comprehensive table. This domain-sharing event table includes the following columns: user_id, domain, age, gender, race, party, party registration, state, and year-month. Each row corresponds to a single sharing event, with users who share the same domains multiple times contributing multiple rows. Due to the presence of user identifiers, we cannot release this detailed table. Instead, we provide aggregate statistics derived from this table including DomainDemo-multivariate and DomainDemo-univariate.
@@
DomainDemo-multivariate describes the domain sharing behavior of users across different demographic dimensions simultaneously. It includes several variants designed to facilitate different types of research analyses. The most granular variant is the monthly distribution data at the domain level, which is produced by grouping the domain-sharing event table by domain, age, gender, race, party (party registration is excluded here), state, and year-month. In each bucket, we calculate the following statistics: the number of shares, the number of unique users who share the domain, and the Gini index of the sharing count across users. Formally, the Gini index $G$ for a domain is calculated as:
$G$ $=\frac{1}{2N^{2}\bar{x}}\sum_{i=1}^{N}\sum_{j=1}^{N}|x_{i}-x_{j}|\text{,}$ (1)
@@
where $N$ is the number of users who share the domain, $x_{\alpha}$ is the number of shares by the $\alpha$-th user, and $\bar{x}=\sum_{i=1}^{N}x_{i}/N$ is the average number of shares per user. $G$ ranges from 0 to 1, with 0 indicating equal sharing and values close to 1 indicating that a few users share the domain disproportionately. Note that we set the Gini index to 1 for domains shared by only one user in the bucket. We include the Gini index to help researchers understand the inequality of sharing events across users without releasing detailed information about these users due to privacy concerns.
In addition to the demographic distribution of users sharing each domain, it is useful to understand the distribution of the whole population in many cases. Therefore, we also include a “baseline” variant that groups the domain-sharing event table by age, gender, race, party, state, and year-month. In each bucket, we calculate the same statistics as the monthly distribution data. We also provide the average number of unique domains shared by users in each demographic bucket and the corresponding standard deviation. On top of the monthly data, we further provide the distribution and baseline data covering the whole time period. In total, DomainDemo-multivariate includes four variants.
@@
DomainDemo-univariate is generated by aggregating the sharing events across all demographic dimensions except for the one of interest. For example, the state univariate data (referred to as the “state universe”) is produced by aggregating the sharing events across all age, gender, race, and party dimensions, resulting in the statistics of sharing events within different states. In each bucket of DomainDemo-univariate, we provide three statistics: the number of shares, the number of unique users who share the domain, and the Gini index of the sharing count across users. And similar to DomainDemo-multivariate, DomainDemo-univariate includes four variants: monthly distribution, monthly baseline, all-time distribution, and all-time baseline.
@@
Domain Localness Metric
Here, we leverage the state universe data from DomainDemo-univariate to derive a data-driven metric that quantifies the “localness” of news domains. This is achieved by calculating the deviation of the user distribution of each domain in different states from the baseline distribution, both of which are provided in DomainDemo-univariate.
For a formal definition, let $C_{s}$ represent the number of unique users in state $s$ across all domains, and $F_{s}$ represent the corresponding frequency, where $F_{s}=C_{s}/\sum_{s}C_{s}$. $F_{s}$ characterizes the baseline distribution of the whole population. For a domain $\delta$, we calculate the user frequency in state $s$, $F_{\delta,s}=C_{\delta,s}/\sum_{s}C_{\delta,s}$, where $C_{\delta,s}$ represents the number of unique users in state $s$ who share the domain $\delta$. For domains shared by diverse users, the observed distribution $F_{\delta,s}$ should closely align with the baseline distribution $F_{s}$ across different states. However, deviations from the baseline distribution are expected for domains with a more concentrated audience.
@@
where $F_{\delta,s}\log_{2}(F_{\delta,s}/F_{s})$ measures the discrepancies between the observed sharing patterns and the baseline distribution of domain $\delta$ in state $s$. $\mathcal{L}_{\delta}$ is a non-negative value that is minimized at zero when $F_{\delta,s}$ and $F_{s}$ are identical. In other words, national news domains should have $\mathcal{L}_{\delta}$ close to zero, while local news domains should have bigger $\mathcal{L}_{\delta}$ values. In the Technical Validation section, we show that $\mathcal{L}_{\delta}$ is a good proxy for the localness of news domains.
@@
A limitation of $\mathcal{L}_{\delta}$ is that it can only indicate the deviation of a domain’s sharing pattern from the baseline distribution. To reveal which states are over-represented or under-represented, one needs to further inspect the values of $F_{\delta,s}$ and $F_{s}$.
Domain Audience Partisanship Metric
Here, we employ the party (and party registration) universe data from DomainDemo-univariate to create data-driven metrics that assess the audience partisanship of domains. We focus on Democrats and Republicans and exclude Independent users. The number of users from each party allows us to quantify the partisanship of the audience for each domain. It is important to note that our audience-based metrics do not evaluate the content characteristics of these domains. However, previous research indicates that audience characteristics are closely associated with the leanings of these domains (Gentzkow and Shapiro, 2010; Buntain et al., 2023).
@@
Formally, the audience partisanship score $\mathcal{P}_{\delta}$ of a domain $\delta$ is calculated as follows:
$\mathcal{P}_{\delta}$ $=\frac{\frac{C_{\delta,r}}{C_{r}}-\frac{C_{\delta,d}}{C_{d}}}{\frac{C_{\delta,r}}{C_{r}}+\frac{C_{\delta,d}}{C_{d}}}\text{,}$ (3)
@@
where $C_{\delta,r}$ and $C_{\delta,d}$ (available in the distribution variant of DomainDemo-univariate) represent the number of unique users from the Republican and Democrat parties who share the domain $\delta$, respectively. $C_{r}$ and $C_{d}$ (available in the baseline variant of DomainDemo-univariate) represent the total number of unique users in the Republican and Democrat parties who share any domain, respectively. Since a user can share multiple domains, we have $C_{\delta,r}\leq C_{r}\leq\sum_{i}C_{i,r}$ and $C_{\delta,d}\leq C_{d}\leq\sum_{i}C_{i,d}$. $\mathcal{P}_{\delta}$ is a continuous value between $-1$ and $+1$, where $-1$ means the domain $\delta$ is exclusively shared by Democratic users and $+1$ means $\delta$ is exclusively shared by Republican users.
Other Metrics
In addition to the localness and audience partisanship metrics, we release three more audience-based metrics: age deviation, race deviation, and gender leaning, to help researchers understand the sharing patterns conditioned on these demographic variables. The age and race deviation metrics are calculated using Eq. (2), where the state categories are replaced with the age or race categories. These metrics quantify how concentrated the audience is in certain age or race groups. The gender leaning metric is calculated using Eq. (3), where the party categories are replaced with the gender categories. Similar to the audience partisanship metric, the gender leaning metric is also a continuous value between $-1$ and $+1$, where $-1$ means the domain is exclusively shared by male users and $+1$ means the domain is exclusively shared by female users.
The calculation of these metrics is very flexible. While we primarily use the unique number of users in both Eq. (2) and Eq. (3), our experiments demonstrate that using the number of shares produces highly correlated results. The metrics can also be calculated over different time periods. In this paper, we present results for the entire time period in the released version, case studies, and validation. To facilitate reproducibility and customization, we provide the code for calculating these metrics, allowing readers to modify the formulas according to their specific needs.
Our formulas in Eq. (2) and Eq. (3) have a limitation: they rely solely on user distribution without accounting for variations in sharing patterns across demographic groups. For example, our analysis reveals that Democratic users share more diverse domains than Republican users, averaging 74.9 unique domains compared to 54.5 across the whole period. To enable researchers to develop more sophisticated metrics that incorporate these behavioral differences, we provide the mean number of unique domains shared by users in each demographic category and the corresponding standard deviations in the baseline variants of our datasets.
Case Studies
Figure 1: Distributions of the derived metrics for all domains, along with detailed information for three domains: cnn.com, news9.com, and wickedlocal.com. The left column presents the joint distributions of our derived metrics and the unique number of users who share each domain across the entire dataset. The color coding represents the number of domains within each grid cell. The symbols indicate the locations of the three domains. The three columns on the right provide detailed distributions of users in various demographic dimensions for the three domains, respectively. Sub-figures (b-d) highlight the discrepancies between the observed user distribution and the baseline distribution across U.S. states for the three domains. The color coding indicates the $F_{\delta,s}\log_{2}(F_{\delta,s}/F_{s})$ value in each state. The bar plots display both the baseline distribution and the distribution of users in each demographic category for the domain of interest. The baseline distribution represents the patterns observed across all domains in the dataset.
To help the readers interpret the derived metrics, we present the distributions for all domains in the dataset and provide case studies for three example domains in Figure 1.
Firstly, cnn.com, a national news outlet, has a user base closely aligned with the baseline. Consequently, its localness ($\mathcal{L}_{\delta}=0.013$), race deviation ($\mathcal{L}_{\delta}=0.002$), age deviation ($\mathcal{L}_{\delta}=0.049$), and gender leaning ($\mathcal{P}_{\delta}=-0.033$) scores are near zero. cnn.com is shared more often by Democratic users and less often by Republican users than the baseline, resulting in an audience partisanship score of $-0.132$.
Due to space constraints, we can only provide three case studies here. We have released an interactive app to allow readers to explore the patterns of other domains in our dataset at domaindemo.info.
@@
Data Access
Our dataset is available on Zenodo (https://doi.org/10.5281/zenodo.15151613(Yang et al., 2025). Given the sensitive nature of the information about Twitter users in our datasets, we have implemented layered access controls. Since DomainDemo-multivariate and DomainDemo-univariate can potentially reveal the identities of Twitter users in the dataset when combined with other datasets, restrictions are imposed on access to them. Specifically, researchers must complete an application process and sign a data use agreement that prohibits the identification of individual Twitter users and re-distribution of the data. Those interested in accessing these datasets can follow the instructions on the Zenodo page. The derived metrics of the domains, such as localness and audience partisanship scores, are made publicly available.
Data Format
Figure 2: Folder structure of the DomainDemo dataset. We release the multivariate and univariate versions of the domain-sharing statistics. We provide both monthly and all-time variants of the data. Each variant contains distribution and baseline subdirectories. For clarity and due to space limitations, we use wildcards to represent patterns in file names rather than listing each file individually. Specifically, {demo} stands for different demographic factors (i.e., age, gender, race, state, party, and partyreg, six universes in total), and {YYYY-MM} denotes year-month combinations from May 2011 to April 2022 (132 months in total). We also release the derived metrics for domains based on the whole time period. The number of files represented by each pattern is indicated with comments on the right.
Figure 2 illustrates the folder structure of the DomainDemo dataset. Due to file count limitations on the data hosting platform, each root folder is distributed as a compressed archive. After downloading and extracting these archives, users will find the subfolders and files organized according to the structure depicted in Figure 2. All data files are provided in CSV format and compressed using the Gzip algorithm for efficient storage and transmission. Users with access can load and analyze them using preferred programming languages, such as Python and R. In the corresponding code repository (see details in the Code Availability section), we provide example scripts to work with the data. In the following, we provide the schema of the tables in DomainDemo.
Domain-sharing Statistics
Table 1: Schema of the DomainDemo-multivariate tables. The “domain” column is only included in the distribution variants, whereas the “domains_count_mean” and “domains_count_std” columns are only included in the baseline variants.
Column Type Category Note
domain string key second-level domains of websites, e.g., nytimes.com; only included in distribution variants, not in baseline variants
time string time year and month in the format “YYYY-MM” for monthly data; equal to “alltime” for all-time data
shares integer statistic sharing event count in the bucket
users integer statistic unique number of users in the bucket
domains_count_mean float statistic average number of unique domains shared by users in the bucket
domains_count_std float statistic standard deviation of the number of unique domains shared by users in the bucket
state string demographic two-letter abbreviations, e.g., MA; including 50 U.S. states and the District of Columbia
@@
The table schema for DomainDemo-multivariate is provided in Table 1. Note that different variants of the dataset have slightly different columns, and party registration is not included. Considering the monthly distribution variant, a row with the following values: domain=example.com, state=CA, race=Asian, gender=Female, age=30-49, party=Democrat, year_month=2018-12, shares=50, users=10, gini=0.1 indicates that the domain example.com was shared 50 times in December 2018 by 10 users who live in California, are Asian females aged between 30 and 49, and identify as Democrats. The Gini index of 0.1 indicates that the sharing count is almost evenly distributed across these users. The baseline data, on the other hand, do not have the domain column. A row with the following values: state=CA, race=Asian, gender=Female, age=30-49, party=Democrat, year_month=2018-12, shares=350, users=120, gini=0.5, domains_count_mean=2.2, domains_count_std=3.5 indicates that there were 350 sharing events in December 2018 by 120 users who live in California, are Asian females aged between 30 and 49, and identify as Democrats. These users shared an average of 2.2 unique domains with a standard deviation of 3.5.
Table 2: Schema of the DomainDemo-univariate tables. There are six universes: state, race, gender, age, party, and party registration. Each universe has a unique column for the corresponding demographic variable but shares other common columns. The “domain” column is only included in the distribution variants, whereas the “domains_count_mean” and “domains_count_std” columns are only included in the baseline variants.
Column Type Category Note
domain string key second-level domains of the websites, e.g., nytimes.com; only included in distribution variants, not in baseline variants
time string time year and month in the format “YYYY-MM” for monthly data; equal to “alltime” for all-time data
shares integer statistic sharing event count in the bucket
users integer statistic unique number of users in the bucket